Write the equation for this reaction, then balance the equation.

B. Calculate the grams of lead(II) iodide that can be produced from 5.00 moles of

potassium iodide.

C. Calculate the grams of lead (II) iodide that can be reproduced from 75.00

grams of potassium iodide


1
Expert's answer
2022-03-23T09:16:04-0400

2KI = PbI2

The mole ratio is 2 : 1, so, moles of PbI2 = 5.00 / 2 = 2.50 mol

C. Calculate the grams of lead (II) iodide that can be reproduced from 75.00

grams of potassium iodide

moles KI = 75.00 / 166 = 0.45 mol

moles PbI2 = 0.45 / 2 = 0.226 mol